# Knowledge Transfer from High-Resource to Low-Resource Programming Languages for Code LLMs

### v5

@@ 5.1. Fine-Tuning Small, Medium, and Large Code LLMs with MultiPL-T
    For Racket and OCaml, which are the lowest-resource languages that we evaluate, the fine-tuned 1B parameter models perform as well as StarCoderBase-15B at baseline.
@@
    Racket and OCaml show the largest relative gains. For example, the fine-tuned versions of both models have more than double the score of their base models, with particularly large gains for OCaml.
@@
    We also fine-tune and evaluate CodeLlama-34B and see significant improvements as well.
Overall, when compared to open models that are not fine-tuned on proprietary data, these fine-tuned models achieve state-of-the-art benchmark scores on Julia, Lua, OCaml, R, and Racket at 1B, 15B, and 34B model sizes.
@@
5.2. Fine-Tuning Efficiency
We now investigate how StarCoderBase-1B’s performance on MultiPL-E varies during fine-tuning. On three languages (Lua, OCaml, and Racket), we fine-tune StarCoderBase-1B for seven epochs of MultiPL-T data and evaluate at each epoch. However, the datasets that we have generated are imbalanced in the number of training items (Table 3). To better balance the training sets, we randomly sample 25,000 items from each dataset to get similarly-sized fine-tuning sets for each language. [^10] However, small differences remain: Lua is more verbose than OCaml, so 25,000 training items for Lua is more data than 25,$0$ training items for OCaml.
@@ 5.2. Fine-Tuning Efficiency
Table 3. Dataset sizes.
@@
Language StarCoderBase-1B StarCoderBase-15B
Base Fine-tuned Base Fine-tuned
OCaml 33.0 33.7 50.6 42.9
Racket 19.3 22.7 28.4 41.3
Lua 26.3 46.9 48.7 51.3
Table 4. The pass@1 scores of the original models and fine-tuned models on our new adversarial benchmark that uses user-defined types, higher-order functions, helper functions, and non-standard libraries. The scores suggest that the 15B model may have overfit when fine-tuned on OCaml. However, the other models do the same or better after fine-tuning.
5.3. A New Adversarial Benchmark
A limitation of the MultiPL-T datasets is that every training item is a single function without any other context: they may use standard libraries, but cannot depend on other functions, classes, or third-party libraries. Thus it is possible that fine-tuning a model on MultiPL-T data will make it overfit to this format. [^11] Recall that the base models have been pretrained on natural code that is not constrained to the MultiPL-T format. Thus appropriate fine-tuning should not eliminate their ability to generate code that is not in the MultiPL-T format. Moreover, conventional Code LLM benchmarks, including MultiPL-E, will not expose this problem, because the benchmark tasks largely involve generating standalone functions that use only standard libraries.
To determine if this kind of overfitting is a problem, we construct a new, multi-language benchmark with fourteen problems. Every problem has hidden test cases and a prompt that exercises the model’s ability to use user-defined types, higher-order functions, helper functions, or external libraries (Table 5). We manually translate these prompts to idiomatic OCaml, Racket, and Lua.
Problem Description
Add Left of NumTree
Context: a type for a tree of numbers. Problem: add the numbers on the left branches.
Add Subway Station
Context: a subway system represented as a graph. Problem: add connections between stations.
Map Over Android Phones
Context: a type for phone models. Problem: a mapping function that only applies to Android phones.
Electrify Instruments
Context: a type hierarchy for several musical instruments with flag that indicates if they are electric. Problem: mark instruments as electric.
Collatz Depth
Context: empty. Problem: requires a solution to see how far a number is from converging on 1 in the Collatz sequence.
Mirror a Tree
Context: a type for trees. Problem: tree reflection.
Double Do It
Context: 2 of 3 helper functions. Problem: define third helper and primary function.
Points and Lines
Context: types for point and lines types Problem: Manhattan distance.
Series
Context: a type which represents a series with its current value and update function. Problem: define a function which updates to the next value in the series.
Shapes
Context: several shape types and functions/methods. Problem: define a function that uses the helpers.
Social Time
Context: types that define calendar events. Problem: calculate time spent on a particular kind of event.
Decode Message
Context: imports a common cryptography library. Problem: Decode AES encrypted message.
Verify Source
Context: imports a common cryptography library. Problem: verify signature.
Start AES
Context: imports a common cryptography library. Problem: generate an AES key and encrypts it.
Table 5. The 14 adversarial problems.
5.4. Evaluating Coding Style
Category Grading Item Max Deduction
Text Dangling parentheses $-0.5$
Line too long $-1$
Using car/cdr $-0.5$
cond with round brackets $-0.5$
Definitions let expression not at beginning of function body $-1$
Nesting define or let expressions $-2$
Unecessary use of let* or letrec $-1$
Defining useless local variables $-1$
Not defining helpers or variables for reused code $-1$
Conditionals Nested if expression instead of cond $-2$
Using (if COND #t #f) $-1$
Traversal Using iteration when recursion is available $-3$
Figure 9. Our Racket style rubric for grading generated programs. The maximum score for a program is 15 points. Items are grouped into categories according to the type of error they entail.
A potential limitation of MultiPL-T is that it may negatively impact the style of generated code, since the training items are translated from Python. We study this issue in Racket using a qualitative evaluation process. We developed a Racket style grading rubric (Figure 9) based on the Racket style guide and our experience teaching and grading Racket programming assignments. The rubric outlines grading items and their corresponding deductions. Several deductions are designed for style problems that arise in code written by beginning students, such as needlessly long lines. Others penalize Lisp style, such as using car and cdr instead of first and rest. Finally, the largest deduction is for using imperative iteration when a simple functional solution is possible.
We use our rubric to grade the HumanEval solutions produced by StarCoderBase-1B before and after fine-tuning on MultiPL-T Racket data. We grade only the 35 problems that both models are able to solve at least once (thus we omit several problems solvable only after fine-tuning). Since we generate several solutions for each problem, we select the most common working solution, and in case of ties, we select the one with the best style. We grade 70 Racket programs in total: 35 produced by the base model and 35 after fine-tuning. To avoid bias, we use two graders and anonymized the selection and grading process by assigning random IDs to each program and tracking their provenance on a hidden spreadsheet. The two graders have substantial Racket teaching experience and we find minimal discrepancy between them: their scores differed by more than 1 point for just 15 out of the 70 candidate programs, and only differed by 3–4 points for two programs.
We inspect the 17 programs that scored higher for the base model than the fine-tuned model. We find that the fine-tuned model is more likely to use nested if expressions in these programs, as well as performing iteration where recursion is available. Conversely, we inspect the eight programs that scored higher in the MultiPL-T model and found that the base model is more likely to use direct recursion with car/cdr while the fine-tuned model uses Racket list abstractions. figure 10 shows the breakdown of the kinds of deductions assigned per problem to each model, where deductions are averaged among graders.
The results of our evaluation are consistent with how we generated our training data from Python code. Overall, although fine-tuning slightly decreases the model’s ability to generate idiomatic Racket code, it also significantly increases its ability to generate correct Racket code as we showed earlier. Moreover, both models perform well on the Racket style rubric, which suggest that the tradeoff is minimal.
Language Basic With Canonical
OCaml 26.1 23.9
Racket 34.7 56.8
Lua 51.4 68.5
5.5. Translation Prompt Ablations
The final part of our evaluation concerns the efficiency of the MultiPL-T pipeline, and not the quality of LLM-generated code. MultiPL-T uses StarCoder-15B to translate training items from Python to low-resource languages. The success rate of this translation is dependent on both the quality of the pretrained model (which is fixed) and the quality of the prompt (which we design). We tried several prompt variations during development, and eventually settled on a prompt that includes the original Python code in a comment (§ 4.3).
Doing a complete ablation with all five languages and ~133,000 functions would be prohibitively expensive. Instead, we do an experiment with a random sample of 1,000 source Python functions. We use the LLM to translate these 1,000 functions to OCaml, Racket, and Lua with two different prompt formats: with and without the original Python source. We generate 50 candidates for each prompt. [^12] Thus this small experiment still requires 300,000 generations from the LLM and takes about 1/2 a day on an A100 GPU. We evaluate performance using pass@50, which is the likelihood that the model produces at least one correct solution in 50 attempts.

### v6

@@ 5.2. Evaluation on Standard Benchmarks
    Compare the new fine-tuned model to the base model. This allows us to ask, how does a fine-tuning approach improve baseline performance for a particular model?
@@
    Compare the new fine-tuned model to other fine-tunes of the same model. Fine-tuning on a large dataset is likely to do better than fine-tuning on a smaller dataset. But, this allows us to ask, How efficient is one fine-tuning approach compared to another?
  3. (3)
    Fine-tune a variety of base models. This allows us to ask, does the fine-tuning approach generalize to different model sizes and families?
Thus a fine-tuning approach should improve baseline performance on several different model families and across several model sizes, and we will show that this holds for MultiPL-T.
Table 2. MultiPL-E pass@1 scores for 1B, 15B, 34B, and 70B parameter models before and after fine-tuned on MultiPL-T data. The fine-tuned models show significant improvement.
Language StarCoderBase-1B StarCoderBase-15B CodeLlama-34B CodeLlama-70B
Base Fine-tuned Base Fine-tuned Base Fine-tuned Base Fine-tuned
OCaml 1.5 9.7 6.9 19.9 18.3 27.4 23.2 29.3
Racket 4.7 11.3 11.8 21.0 15.9 29.1 21.9 33.1
R 5.4 8.9 10.2 17.3 18.2 25.5 23.0 28.5
Julia 11.3 15.6 21.1 35.2 31.8 43.5 41.9 44.5
Lua 12.1 17.3 26.6 31.0 38.1 43.9 41.7 44.9
5.2.1. Fine-Tuning StarCoderBase and Code Llama
Fine-tuning StarCoderBase and Code Llama with MultiPL-T-generated data on our target languages improves performance on MultiPL-E across the board (Table 2). For each model, we fine-tune a separate model for Julia, Lua, OCaml, R, and Racket. We checkpoint and evaluate the models at each epoch and report the peak performance. It is not a goal of this paper to maximize MultiPL-E scores. In fact, the next section suggests that it would be easy to improve some of these scores by either training longer on the data we already have or by letting MultiPL-T generate more data. Beyond the generally improved performance, we draw several other conclusions:
  1. (1)
    Racket and OCaml, which are the lowest-resource languages that we evaluate, show the largest relative gains. For example, the fine-tuned versions of both models have more than double the score of their base models, with particularly large gains for OCaml. Even the fine-tuned 1B models perform comparably to the base 15B models at producing correct solutions in these languages.
@@ 5.2.1. Fine-Tuning StarCoderBase and Code Llama
    We also fine-tune and evaluate CodeLlama-34B and CodeLlama-70B, where we see significant improvements as well.
  5. (5)
    In contrast to distillation, where a larger model produces training data for a small model, the MultiPL-T approach allows a smaller model to improve the performance of much larger models. Specifically, we use data generated by StarCoderBase-15B and validated using the MultiPL-T approach to improve the performance of models that are nearly 5x larger.
The supplementary material includes an evaluation with the MBPP benchmark. On MBPP, we show even larger relative gains with MultiPL-T.
Table 3. MultiPL-E pass@1 scores for two recently released models on Racket. The fine-tuned models are significantly better, even though the MultiPL-T datasets are from an older model.
Language StarCoder2-15B DeepSeekCoder-Base-33B
Base Fine-tuned Base Fine-tuned
Racket 22.41 29.7 23.3 36.6
5.2.2. Fine-Tuning StarCoder2 and DeepSeek Coder
When this work was done, Code Llama was the best-performing base Code LLM with open model weights. However, two new base models were recently released: DeepSeek Coder (Guo et al., 2024) and StarCoder2 (Lozhkov et al., 2024). To show that MultiPL-T remains relevant, we fine-tune DeepSeekCoder-33B and StarCoder2-15B on our Racket dataset. The aforementioned models are the two top-performing base models of their size at the time this article was written. In both cases, we find that the MultiPL-T approach continues to work. Both models show significant improvement in their Racket capabilities (table 3).
The datasets presented in this article are part of the StarCoder2 pretraining corpus as one of several sources of “(Leandro’s) High Quality” data. All we can conclude is that they did not hurt performance on the target languages and that further fine-tuning still help (contrast with § 3.1).
5.2.3. Summary
Using MultiPL-T, we have fine-tuned models from four different model families and a wide range of model sizes (1B to 70B parameters). We find that our fine-tuned models outperform base models across the board, with the most significant gains for the lowest-resource languages and small to mid sized models.
The largest models (70B) do not show as much as the smallest models. However, our datasets are generated with a much smaller model (15B), and it is unusual that we can use a smaller model to improve the performance of a much larger model (Burns et al., 2024).
At the time of writing, our fine-tuned version of DeepSeek Coder outperforms every fine-tuned model on the BigCode Models Leaderboard for Racket (Allal, 2024). The leaderboard models fine-tuned on an order of magnitude more data than ours, as well as models that distill proprietary models such as GPT-4. Thus we conclude that for low-resource languages, MultiPL-T is significantly more data efficient than alternative approaches to fine-tuning.
Table 4. The pass@1 scores of the original models and fine-tuned models on our new benchmark that uses user-defined types, higher-order functions, helper functions, and non-standard libraries. The scores suggest that the 15B model may have overfit when fine-tuned on OCaml. However, the other models do the same or better after fine-tuning.
Language StarCoderBase-1B StarCoderBase-15B
Base Fine-tuned Base Fine-tuned
OCaml 33.0 33.7 50.6 42.9
Racket 19.3 22.7 28.4 41.3
Lua 26.3 46.9 48.7 51.3
5.3. New Tasks and Qualitative Evaluation
The previous section uses MultiPL-E for evaluation, which has a very particular format (§ 2.2) and uses unit tests to test correctness. Code LLMs are much more flexible and there are acceptability criteria that cannot be captured with unit tests or specific formats. We address these in this section.
5.3.1. Evaluating In-Context Learning
A limitation of the MultiPL-T datasets is that every training item is a single function without any other context: they may use standard libraries, but cannot depend on other functions, classes, or third-party libraries. Thus it is plausible that fine-tuning a model on MultiPL-T data will make it overfit to this format. [^9] Recall that the base models have been pretrained on natural code that is not constrained to the MultiPL-T format. Thus appropriate fine-tuning should not eliminate their ability to generate code that is not in the MultiPL-T format. Moreover, conventional Code LLM benchmarks, including MultiPL-E, will not expose this problem, because the benchmark tasks largely involve generating standalone functions that use only standard libraries.
To determine if this kind of overfitting is a problem, we construct a new, multi-language benchmark with fourteen problems. We decided to manually construct each problem as opposed to sourcing them from repositories to ensure that the problems were similar across languages, and more importantly, that they were not part of the training data. Every problem has hidden test cases and a prompt that exercises the model’s ability to use user-defined types, higher-order functions, helper functions, or external libraries (Table 5). We manually translate these prompts into idiomatic OCaml, Racket, and Lua.
Table 5. The 14 problems that exercise in-context learning.
Problem Description
Add Left of NumTree Context: a type for a tree of numbers. Problem: add the numbers on the left branches.
Add Subway Station Context: a subway system represented as a graph. Problem: add connections between stations.
Map Over Android Phones Context: a type for phone models. Problem: a mapping function that only applies to Android phones.
Electrify Instruments Context: a type hierarchy for several musical instruments with a flag that indicates if they are electric. Problem: mark instruments as electric.
Collatz Depth Context: empty. Problem: requires a solution to see how far the given number is from converging on 1 in the Collatz sequence.
Mirror a Tree Context: a type for trees. Problem: tree reflection.
Double Do It Context: 2 of 3 helper functions. Problem: define the third helper and the primary function.
Points and Lines Context: types for point and lines types Problem: Manhattan distance.
Series Context: a type that represents a series with its current value and update function. Problem: define a function that updates to the next value in the series.
Shapes Context: several shape types and functions/methods. Problem: define a function that uses the helpers.
Social Time Context: types that define calendar events. Problem: calculate time spent on a particular kind of event.
Decode Message Context: imports a common cryptography library. Problem: Decode AES encrypted message.
Verify Source Context: imports a common cryptography library. Problem: verify signature.
Start AES Context: imports a common cryptography library. Problem: generate an AES key and encrypts it.
@@ 5.3.1. Evaluating In-Context Learning
[…]
5.3.2. Evaluating Coding Style
Table 6. Our Racket style rubric for grading generated programs. The maximum score for a program is 15 points. Items are grouped into categories according to the type of error they entail.
Category Grading Item Max Deduction
Text Dangling parentheses $-0.5$
Line is too long $-1$
Using car/cdr $-0.5$
cond with round brackets $-0.5$
Definitions let expression not at beginning of function body $-1$
Nesting define or let expressions $-2$
Unnecessary use of let* or letrec $-1$
Defining useless local variables $-1$
Not defining helpers or variables for reused code $-1$
Conditionals Nested if expression instead of cond $-2$
Using (if COND #t #f) $-1$
Traversal Using iteration when recursion is available $-3$
[…]
(define (rolling_max numbers)
(define (max_in_list list)
(foldl max 0 list))
(define (max_in_list_with_index list)
#| body omitted |#)
(let loop ((numbers numbers) (max_numbers ’()))
(if (empty? numbers)
max_numbers
(let* ((number (car numbers)) (max_number (max_in_list max_numbers)))
(loop (cdr numbers)
(append max_numbers
(list (if (> number max_number) number max_number))))))))|#
(a) A solution from the fine-tuned model. The max_in_list_with_index function is redundant.
(define (rolling_max numbers)
(define (max_list numbers)
(if (null? numbers)
(cons (apply max numbers) (max_list (cdr numbers)))))
(reverse (max_list (reverse numbers))))
(b) A better solution from the base model.
Figure 12. Two working solutions to the same HumanEval-Racket problem produced by the base and MultiPL-T fine-tuned model. In this case, the base model produces higher-quality code.
[…]
A potential limitation of MultiPL-T is that it may negatively impact the style of generated code, since the training items are translated from Python. We study this issue in Racket using a qualitative evaluation process. We developed a Racket style grading rubric (Table 6) based on the Racket style guide and our experience teaching and grading Racket programming assignments. The rubric outlines grading items and their corresponding deductions. Several deductions are designed for style problems that arise in code written by beginning students, such as needlessly long lines. Others penalize Lisp style, such as using car and cdr instead of first and rest. Finally, the largest deduction is for using imperative iteration when a simple functional solution is possible.
We use our rubric to grade the HumanEval solutions produced by StarCoderBase-1B before and after fine-tuning on MultiPL-T Racket data. We grade only the 35 problems that both models can solve at least once (thus we omit several problems solvable only after fine-tuning). Since we generate several solutions for each problem, we select the most common working solution, and in the case of ties, we select the one with the best style. We grade 70 Racket programs in total: 35 produced by the base model and 35 after fine-tuning. To avoid bias, we use two graders and anonymized the selection and grading process by assigning random IDs to each program and tracking their provenance on a hidden spreadsheet. The two graders have substantial Racket teaching experience and we find minimal discrepancy between them: their scores differed by more than 1 point for just 15 out of the 70 candidate programs, and only differed by 3–4 points for two programs.
We inspect the 17 programs that scored higher for the base model than the fine-tuned model. We find that the fine-tuned model is more likely to use nested if expressions in these programs, as well as performing iteration where recursion is available. Conversely, we inspect the eight programs that scored higher in the MultiPL-T model and find that the base model is more likely to use direct recursion with car/cdr while the fine-tuned model uses Racket list abstractions. figure 11 shows the breakdown of the kinds of deductions assigned per problem to each model, where deductions are averaged among graders. Figure 12 shows a graded pair where the base model scores better on coding style than the fine-tuned model, and the supplementary material shows a graded pair where the converse is true: the fine-tuned model produces a cleaner solution than the base model. Coding style aside, in both cases the lower quality solution is much more verbose than necessary.
The results of our evaluation are consistent with how we generated our training data from Python code. Overall, although fine-tuning slightly decreases the model’s ability to generate idiomatic Racket code, it also significantly increases its ability to generate correct Racket code as we showed earlier. Moreover, both models perform well on the Racket style rubric, which suggests that the trade-off is minimal.
Language Basic With Python Solution
OCaml 26.1 23.9
Racket 34.7 56.8
Lua 51.4 68.5
5.4. MultiPL-T Ablations
The final part of our evaluation are ablations to demonstrate the efficacy of different parts of the MultiPL-T pipeline. This section showcases results for a significant number of fine-tuned models. To manage costs, we use StarCoderBase-1B for the majority of the experiments.
5.4.1. Fine-Tuning Efficiency
Table 8. Dataset sizes.
[…]
@@ 5.4.1. Fine-Tuning Efficiency
We now investigate how the performance of StarCoderBase-1B on MultiPL-E varies during fine-tuning. On three languages (Lua, OCaml, and Racket), we fine-tune StarCoderBase-1B for seven epochs of MultiPL-T data and evaluate at each epoch. However, the datasets that we have generated are imbalanced in the number of training items (Table 8). To better balance the training sets, we randomly sample 25,000 items from each dataset to get similarly-sized fine-tuning sets for each language. [^10] However, small differences remain: Lua is more verbose than OCaml, so 25,000 training items for Lua is more data than 25,000 training items for OCaml.
@@
5.4.2. Translation Versus Generation
MultiPL-T uses StarCoderBase-15B to translate training items from Python to low-resource languages. The success rate of this translation is dependent on both the quality of the pretrained model (which is fixed) and the quality of the prompt (which we design). We tried several prompt variations during development and eventually settled on a prompt that includes the original Python code in a comment (§ 4.3).
Doing a complete ablation with all five languages and ~133,000 functions would be prohibitively expensive. Instead, we run an experiment with a random sample of 1,000 source Python functions. We use the LLM to translate these 1,000 functions to OCaml, Racket, and Lua with two different prompt formats: with and without the original Python source. We generate 50 candidates for each prompt. [^11] Thus this small experiment still requires 300,000 generations from the LLM and takes about 1/2 a day on an A100 GPU. We evaluate performance using pass@50, which is the likelihood that the model produces at least one correct solution in 50 attempts.
@@ 5.4.2. Translation Versus Generation
Table 9. Without test-based validation, the LLM-translated training items are poor quality and fine-tuning is not effective. The performance of the 15B model decreases slightly compared to the base model. The performance of the 1B model increases slightly, but is far worse than the result we get with the tested items.
Model Baseline MultiPL-T No Validation
StarCoderBase-1B 4.7 11.3 8.6
StarCoderBase-15B 11.8 21.0 11.6
5.4.3. Testing Is Critical for MultiPL-T
MultiPL-T relies on tests to validate the LLM-translated training items. We reason that since we are translating to programming languages on which the LLM performs poorly at synthesis, it is also likely to perform poorly at translation. But, to validate this claim, we fine-tune models on Racket data without test validation. We get poor results as expected (table 9). This shows that the effort we take to generate, validate, and compile tests is essential for MultiPL-T.
5.4.4. The Impact of Deduplication
Prior work has shown that data duplication decreases performance while increasing training time (Lee et al., 2022; Allal et al., 2023). To demonstrate its impact in our work, we do an experiment on the dataset where it has the most impact—where deduplication discards the most data—which is for Lua. This is to be expected because the number of nearly duplicate functions will be higher when the data generating LLM translates functions with a higher success rate. Of the languages we target, our data generating LLM (StarCoderBase15-B) has the highest pass rate on Lua.
Before deduplication, the Lua dataset has 1.4M functions, but after deduplication we are left with 48K (reported in table 8). For this experiment, we fine-tune StarCoderBase-1B on a single H100 GPU on the undeduplicated dataset, with exactly the same hyperparameters we used earlier (§ 5). We find the following:
  1. (1)
    Without deduplication, the fine-tuned StarCoderBase-1B gets a 17.1 pass@1 score on MultiPL-E and takes 12 hours to train.
  2. (2)
    With deduplication, the fine-tuned StarCoderBase-1B gets 17.3 pass@1 on MultiPL-E and takes less than 30 minutes to train.
The two pass@1 scores are very close. But, without deduplication, training takes significantly longer, because of the increase in dataset size.
The impact would be more significant with a larger model. For example, it took us 8 hours on 8xH100 GPUs to fine-tune StarCoderBase-15B with a deduplicated dataset. Without deduplication, fine-tuning would take several days. Similarly, fine-tuning the 33B and 70B models would take even longer. Thus deduplication is a way to manage the time and cost of training large models.
5.4.5. Targeting the Source Language
The complete MultiPL-T toolchain is designed to produce training data for a target low-resource language from a source language (Python). However, it is also possible to directly evaluate the first part of the MultiPL-T toolchain, which builds a high-quality dataset for the source language (Python). In this section, we fine-tune select LLMs on Python using the dataset of 133,668 Python functions that are documented, tested, typeable, and filtered on test coverage (§§ 4.1 and 4.2).
We fine-tune and evaluate StarCoderBase-1B directly on this dataset of Python functions. The base model gets a 15.1 pass@1 score on Python, whereas the fine-tuned model gets 15.0 pass@1.121212We observe this score on epoch 3. On epochs 1–7, the score varies between 10.7 and 15.0. Thus, it is clear that fine-tuning directly on this subset of Python does not produce the same improvements that we see on the low-resource languages after translation.
We hypothesize that this occurs for the following reason: StarCoderBase-1B has already been trained on three epochs of this Python data. We are not fine-tuning the model on novel data, thus there is little new for the model to learn, and this is reflected in the unchanged pass@1 score. In contrast, when we translate this data to a low-resource language using the full MultiPL-T pipeline, we are effectively creating novel data for the model to learn.